1) Calculate the number of moles of atoms contained in 2.35grams Lithium?
1
Expert's answer
2020-09-17T07:18:35-0400

Solution:

Mole: the amount of a substance that contains 6.02×1023 respective particles of that substance.


Moles of lithium = Mass of lithium / Molar mass of lithium

The molar mass of lithium is 6.941 g/mol.

Thus:

Moles of lithium = (2.35 g) / (6.941 g/mol) = 0.338 mol = 0.34 mol


The number of atoms of lithium = Moles × Na = 0.34 mol × 6.02×1023 = 2.047×1023 atoms of lithium


Answer: 0.34 mol is the number of moles of lithium.

The number of atoms of lithium is 2.047×1023 atoms.